OpenStudy (awy):

Find the p-value given the following null and alternative hypothesis with a test statistics of 2.153. Null: p=0.56 Alternative: p not equal 0.56

rvc (rvc):

it is binomial distribution?

OpenStudy (awy):

i dont think so, this is it asked

OpenStudy (awy):

and given

OpenStudy (anonymous):

It looks like \(p=0.56\) is a proportion as part of the null hypothesis. You want to find the estimated proportion \(\hat{p}\) given that the test statistic \(Z=2.153\). The \(Z\) statistic for proportions is given by \[Z=\frac{\hat{p}-p}{\sigma_p}=\frac{\hat{p}-p}{\sqrt{\dfrac{p(1-p)}{n}}}\] Plug in everything you know: \[2.153=\frac{\hat{p}-0.56}{\sqrt{\dfrac{0.56(1-0.56)}{n}}}\] Ideally, you would solve for \(\hat{p}\), but without a sample size I'm afraid there's not much you can do in the way of finding an actual value for \(\hat{p}\)...

OpenStudy (awy):

Alright, thx for the reply. This was a exam question and quite a curve ball since the homework and group assignments never included how to find the p-value without a sample size.

OpenStudy (anonymous):

@awy It's possible that I misinterpreted what the question meant by "p-value." This could be the p-value you would obtain from the actual statistical test for significance. However, I don't see how you'd be able to find that either, since you need the significance level of the test. You can probably assume a 5% significance level (that's the standard level) and determine the rejection region for the test. In this case, if we assume a 5% significance level, then the rejection region is \(|Z|>1.96\) (keep in mind this is a two-tailed test). Since \(2.153>1.96\), you reject the null hypothesis. I'm getting that the p-value is 0.031319.